JOHN SYAYS TO BRENDA, "IF YOU GIVE ME SEVEN DOLLARS, WE WILL HAVE AN EQUAL AMOUNT OF MONEY." BRENDA RESPONDS, "THIS MAY BE TRUE, BUT IF YOU I GIVE ME SEVEN DOLLARS, I WILL BE ONE DOLLAR SHY OF HAVING TWICE AS MUCH MONEY AS YOU." HOW MUCH MONEY DO JOHN AND BRENDA EACH HAVE?
Answer by mananth(12269)

AMOUNT OF MONEY." BRENDA RESPONDS, "THIS MAY BE TRUE, BUT IF YOU I GIVE ME SEVEN DOLLARS, I WILL BE ONE DOLLAR SHY OF HAVING TWICE AS MUCH MONEY AS YOU." HOW MUCH MONEY DO JOHN AND BRENDA EACH HAVE?
..
brenda = x
John = y
..
y+7=x-7
x-y=14..............1
...
x+7=2y-1
x-2y=-7-1
x-2y=-8............2
..
multiply equation 1 by -2
-2x+2y=-28
add this to 2
-x=-36
/-1
x=36 Brenda
...
x-y=14
so y= 22 John
...